The ratio of forces between two small spheres with constant charge in air to that in a medium of dielectric constant K?

Grade:10

2 Answers

Vikas TU
7 years ago
Force due to charge in air = > q1q2/(4*pi*e)r^2 = F1
Force due to charge in di – electric medium = F2
or
F2 = q1q2/(4*pi**e*K)r^2
where K is the dielectric constant.
Dividing the forces we get,
F1/ F2 = K
 
ankit singh
3 years ago
 
Force due to charge in air = > q1q2/(4*pi*e)r^2 = F1
Force due to charge in di – electric medium = F2
or
F2 = q1q2/(4*pi**e*K)r^2
where K is the dielectric constant.
Dividing the forces we get,
F1/ F2 = K